Rhona has apples that sell 80 pesos and 100 pesos per kilogram. How many kilograms of each must be mixed together to make 100 kilograms of apple that would be sold at 82 pesos per kilogram?
Answer by stanbon(75887)

Rhona has apples that sell 80 pesos and 100 pesos per kilogram. How many kilograms of each must be mixed together to make 100 kilograms of apple that would be sold at 82 pesos per kilogram?
---------
Equations::
Quantity:: e + t = 100 kg
Value:::: 80e+100t = 82*100 pesos
----------------------------------------
Modify for elimination:
8e + 8t = 8*100
8e + 10t = 8.2*100
-------------------
Subtract and solve for "t":
2t = 0.2*100
t = 10 kg (amt of 100 peso apples needed)
---
e = 100 - 10 = 90 kg (amt of 80 peso apples needed)
-----------
Cheers,
Stan H.
-----------------